OpenStudy (anonymous):

You are selling tickets for a high school basketball game. Student tickets cost $3 and general admission tickets cost $5. You sell 350 tickets and collect $1,450. How many student tickets did you sell? Hey do you also know how to do this?

OpenStudy (anonymous):

@abriannaridley

OpenStudy (anonymous):

Ok ready?

OpenStudy (anonymous):

Im ready lol

OpenStudy (anonymous):

ok this is also going to be a system. let x = student ticket and y = general admissions x + y = ?

OpenStudy (anonymous):

actually x = # of student ticket and y = # of general admissions

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Okay got it

OpenStudy (anonymous):

x+y = 350 3x + 5y = 1450 Now we need to solve the system.

OpenStudy (anonymous):

remember how?

OpenStudy (anonymous):

Uhmm we have to solve for Y? my memory is sucky :P

OpenStudy (anonymous):

haha ok: x+y = 350, solve it for y y = 350 - x Then plug into the second equation and solve for x value: 3x + 5(350 -x) = 1450 3x + 1750 - 5x = 1450 -2x + 1750 = 1450 -1750 on both sides -2x = -300 /-2 on both sides x = 150 Now plug in to first equation: 150 + y = 350 y = 200

OpenStudy (anonymous):

Therefore you sold 150 student tickets. <-- Final answer

OpenStudy (anonymous):

Sorry that took so long... Got it?

OpenStudy (anonymous):

OMG YOUR A LIFE SAVER!!! and yes i got it this time. My school feels no need to explain anything and they just expect you to know it -.-

OpenStudy (anonymous):

Mine too, I see a trend... I got AP Chem Home work today without the teacher teaching anything. It is pretty hard

OpenStudy (anonymous):

ahah and im over here complaining about Algebra I lol im lame :P

OpenStudy (anonymous):

Never - I love these questions

OpenStudy (anonymous):

i'll keep that in mind lol i dont understand it at all!